OpenStudy (anonymous):

Do i sudtract or Divide or Multiply?? 21 meter by 0.26

OpenStudy (anonymous):

please include the full question - what you wrote doesn't quite make sense.

OpenStudy (anonymous):

Loren is putting bricks along both edges of the 21-meter walkway to his house. Each brick is 0.26 meters long. Loren is placing bricks end to end. How many bricks does be need?

OpenStudy (anonymous):

To get the concept, think easier numbers first. Pretend the walk is 5 feet, and each brick is 1 foot. How many bricks would he need?

OpenStudy (anonymous):

"divide" is right, but you probably could answer your own question by thinking about it with easy numbers

OpenStudy (anonymous):

5 feet

OpenStudy (anonymous):

5 bricks, you mean... :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So it is dividing up the length into brick-sized sections... the number of sections is the number of bricks he needs. So in the easy example, if it was 10 feet long, each brick is 1 ft still, you need 10/1 = 10 bricks.

OpenStudy (anonymous):

In the real problem, you need (21 m long sidewalk)/(0.26 m long brick) = number of bricks

OpenStudy (anonymous):

Does he have 1.2385 bricks??

OpenStudy (anonymous):

@emelylove sorry, I was away a few minutes... i don't think you did the division right. I got 21/0.26 = 80.77 bricks.